# On Generating Monolithic and Model Reconciling Explanations in Probabilistic Scenarios

### v1

@@ 5       $I_{k}\leftarrow\texttt{getIntersections}(\mathcal{B}^{\downarrow w},\Omega_{\varphi},k)$
6       if not SAT$(I_{k}\cup\{\lnot\varphi\})$ then
7             ϵ~~italic-$\tilde{\epsilon}\leftarrow$ monolithic-explanation$(I_{k},\varphi)$
@@ 11             $k\leftarrow k-1$
Algorithm 3 probabilistic-monolithic-explanation$(\mathcal{B},\varphi,\hat{k})$
@@
The proof rests on the fact that, in the worst case, the parameter $k$ will reach a value of $1$. This will then correspond to the most-probable world of $\varphi$, which entails all possible monolithic explanations for $\varphi$. From Lemma 2, we know that the most-probable world of $\varphi$ entails all possible monolithic explanations for $\varphi$, that is, for any $\tilde{E}(\varphi)$, $\omega_{1}\models\tilde{\epsilon}$, and $\tilde{\epsilon}\subseteq I_{1}$. Therefore, as Algorithm 3 uses $I_{1}$ as an input to Algorithm 1, which is guaranteed to return a solution, the algorithm is also guaranteed to terminate with a solution.
@@
[…]
@@ 4.2.2 Model Reconciling Explanations
We now move on to the case of computing probabilistic model reconciling explanations ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for an explanandum $\varphi$ from an agent knowledge base $\mathrm{KB}_{\alpha}$ for a human belief base $\mathcal{B}_{h}$. Similarly to what we described in Section 4.2.1, Algorithm 2 can directly be used on $\mathrm{KB}_{\alpha}$ and $\mathcal{B}_{h}^{\downarrow w}$ for computing model reconciling explanations. Additionally, the concept of a $k$-bounded explanation (Definition 19) can also be used to guarantee a lower bound on the probability of $\tilde{\epsilon}^{+}$.
@@
Result: A probabilistic model reconciling explanation ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$
@@ 14             $k\leftarrow k-1$
Algorithm 4 probabilistic-model-reconciling-explanation$(\mathrm{KB}_{\alpha},\mathcal{B}_{h},\varphi,\hat{k})$
@@ 5.1 Experimental Setup
Experiments were conducted on a system equipped with an M1 Max processor and 32GB of memory. The algorithms were implemented in Python, utilizing the PySAT toolkit (?) for SAT solving, MCS/MUS finding, weighted MaxSAT, and minimal hitting set computations. The time limit for all experiments was set to $500s$.
@@
Note that we created associated belief bases for each problem by simply adding a random weight to each formula in the knowledge base. We will make the code and all benchmarks publicly available in the final version of the paper.
@@ 5.2.1 Monolithic Explanations
Table 3: Number of Instances Solved (S) vs. Timed Out (T/O) by alg1 ($\hat{k}=1$) and alg3 ($\hat{k}=50$, $\hat{k}=100$, $\hat{k}=150$, $\hat{k}=200$).
@@
Table 3 tabulates the instances solved (i.e., found a monolithic explanation within the time limit) and not solved (i.e., timed out) by alg1 ($\hat{k}=1$) and alg3 at $\hat{k}=\{5,100,150,200\}$.151515alg3 at $\hat{k}=1$ corresponds to alg1 because each encoded knowledge base is consistent and entails the explanandum. As such, all formulae in the knowledge base are true in the most-probable world of the explanandum (i.e., $\hat{k}=1$), which means that alg3 reduces to alg1. We observe that the algorithm managed to solve most instances across different values of $\hat{k}$. Figure 1 shows the runtime distributions of alg3 across all values of $\hat{k}$ for computing a monolithic explanation. Interestingly, we observe that the runtimes decrease as $\hat{k}$ increases. This can be explained by the fact that for larger values of $\hat{k}$, alg3 considers the intersections of more worlds where the explanandum is true, which means that the number of formulae that are true in these intersections decreases. As such, the overall search space of monolithic explanations decreases as well, thus resulting in a reduced runtime needed for alg1 to compute a monolithic explanation. This can also be observed more granularly in Figure 2, where we can see the runtime distributions of alg1 ($\hat{k}=1$) and alg3 at $\hat{k}=200$ for each instance of the planning, scheduling, and random CNF problems. Again, the runtime of alg3 at $\hat{k}=200$ is smaller than that of alg1. Moreover, and as expected, in Figure 3, we can observe a positive correlation between runtime and the size of the encoded knowledge bases – as the size of the knowledge base increases, the runtimes increase as well. This is due to the fact that there is an increasing number of variables and formulae that must be considered, thus increasing the computational effort needed by the WMaxSAT, MCS, and hitting set solvers.
@@ 5.2.2 Model Reconciling Explanations
Table 4: Instances Solved (S) vs. Timed Out (T/O) for the Planning, Scheduling, and Random CNF Benchmarks for alg2 and alg4 at $\hat{k}=200$.
@@
Table 4 tabulates the instances solved and timed out by alg2 and alg4 at $\hat{k}=200$ across the five scenarios, where we observe the following trends. For the planning instances, the runtime of both algorithms increases as the difference between the models of the agent and human increases (Scenarios 1 to 5), since both algorithms search over the explanation search space, which increases as the number of differences between the two models increases. As in the previous experiments, alg4 at $\hat{k}=200$ yields faster runtimes than alg2. For the scheduling instances, we observe that the runtimes increase from Scenario 1 to 3, but decrease from Scenarios 4 to 5. Upon closer inspection, this is mainly because the instances solved in these scenarios were easier (i.e., smaller knowledge base sizes) than those solved in the other three scenarios, thus resulting in smaller average runtimes. A similar trend is observed for the random CNF instances. However, in the random CNF instances, alg2 managed to solve more instances than alg4. After examining them more closely, we found that the main bottleneck of alg4 in those instances was computing the most-probable worlds of the explanandum (i.e., the WMaxSAT solver). Even for smaller values of $\hat{k}$, the solver failed to compute all the worlds under the specified time limit – the increase in search space (e.g., because of considering $\mathcal{B}_{\alpha}$ and $\mathcal{B}_{h}$) increased the complexity of these instances. We expect that an optimized and more dedicated solver may be able to overcome this limitation. The runtime distributions for alg2 and alg4 across all commonly solved instances and across commonly solved instances in each scenario can be seen in Figures 4 and 5, respectively. For these instances, we observe, like in the previous experiments, that alg4 has faster runtimes than alg2.
Moreover, in Figure 6 we see the distributions of the model reconciling explanation lengths computed by both algorithms. As expected, the general trend is that the size of the explanation $\epsilon^{+}$ (i.e., formulae from $\mathrm{KB}_{\alpha}$ for $\mathrm{KB}_{h}$ (or $\mathcal{B}_{h}$)) increases with each scenario, as the difference between the agent and human models increase. The same trend can be seen for $\epsilon^{-}$ – each scenario from 1 to 5 has an increasing amount of inconsistencies between the two models. Interestingly, $\epsilon^{-}$ was largest in the random CNF instances. This indicates that the inconsistencies between the human model and the corresponding $\epsilon^{+}$ were high. That can also be used to explain why alg4 failed to solve a subset of random CNF instances – highly inconsistent knowledge bases are considered as the most difficult instances for MaxSAT solvers.
@@ 6.1 Related Planning Work
We have briefly discussed the body of work within explainable AI planning (XAIP) on the topic of model reconciliation in Section 1 as well as situated our work within that literature. We will now discuss other XAIP work, such as contrastive explanations, which take the form of “Why not A (instead of B)?”, where A is an alternative (or foil) suggested by the human to a decision B proposed by the agent. Contrastive explanations have found applications in linear temporal logic systems (?), general epistemic accounts (?), multi-agent optimization problems (?), and in oversubscription planning (?). There have also been used to provide a taxonomy of user questions that often arise during interactive plan exploration (?), as well as towards creating user interfaces for decision support systems (?, ?). For a more thorough discussion of the related work from the XAIP literature, please refer to the discussion by ? (?).
@@ 6.2 Related KR Work
In this paper, we presented a framework for probabilistic explanation generation in monolithic and model reconciliation scenarios. In the monolithic case, our definition of a probabilistic monolithic explanation (Definition 9) may appear similar to what was proposed by ? (?). Nonetheless, an important distinction here is that ? is dealing with epistemic states that do not contain the explanandum, while we are dealing with belief bases that do contain the explanandum. We also define a different notion of explanatory power as well as present algorithms for computing explanations. ? (?) have also considered the problem of defining what constitutes a (monolithic) explanation in probabilistic system, however they focus on epistemic states defined over causal structures.
@@
The notion of (monolithic) explanation has also been explored by the probabilistic logic programming (PLP) community (?, ?), a formalism that extends logic programming languages (i.e., Prolog) with probabilities. In PLP, explanations have been associated with possible worlds. For instance, the MPE (most probable explanation) task consists in finding the world with the highest probability given some evidence (?). However, a world does not show the chain of inferences of a given explanandum and, moreover, it is not minimal by definition, since it usually includes a (possibly large) number of probabilistic facts whose truth value is irrelevant for the explanandum. Another alternative consists in using the proof of an explanandum as an explanation (?), where one can associate a proof with a (minimal) partial world $\omega^{\prime}$ such that for all worlds $\omega\supseteq\omega^{\prime}$, the explanandum is true in $\omega$. In this case, one can easily ensure minimality, but even if the partial world contains no irrelevant facts, it is still not easy to determine the chain of inferences behind a given explanandum. ? (?) have tackled explanation generation in PLP from the perspective of weighted model counting and knowledge compilation.
@@
In the model reconciliation setting, we have extended our previous work on the logic-based model reconciliation problem (?) to a probabilistic case (Definition 14) for capturing scenarios where the human model is uncertain. ? (?) proposed a method for generating explanations in the case of uncertain human models, however, their approach is limited to planning problems, and importantly, it does not quantify the uncertainty levels of the generated explanations, that is, there is no notion of probabilistic explanation. In contrast, the application of probabilistic explanations in the context of model reconciliation that we consider in this work is, to our knowledge, novel.
@@
Finally, the algorithms presented in this paper are an extension of our previous work (?). Specifically, Algorithms 1 and 2 are inspired by a procedure for computing a smallest minimal unsatisfiable set (SMUS) of an inconsistent formula, originally presented by ? (?). The method is also related to other similar approaches for enumerating MUSes and minimal correction sets (MCSes). Moreover, our approach is similar in spirit to the HS-tree presented by ? (?). Although the original purpose was to enumerate diagnoses, Reiter’s procedure can be easily adapted to enumerate MUSes (called conflicts in that paper) as already noted by ? (?). However, the computation of an SMUS might require more substantial modifications. Procedures like the one presented by Reiter, which target MCSes (diagnoses) instead of MUSes (conflicts), can be seen as the dual version of our algorithm. In particular, the algorithm MaxHS (?) applies the same idea of iteratively computing and testing a minimal hitting set for the computation of a MaxSAT solution (the complement of the smallest MCSes). Finally, there are other approaches that exploit the duality between MUSes and MCSes, but instead of iteratively checking if the current hitting set is an MUS, they first compute the set of all MCSes (?). This has the potential advantage that once all the MCSes are known, every minimal hitting set on the collection of all MCSes is guaranteed to be an MUS (Proposition 1). However, as the number of MCSes is, in the worst case, exponential in the size of the formula, this approach might fail even before reporting the first MUS. This is particularly unnecessary when the target is to return a single support, like the one presented in this paper.

### v2

@@ 5 $I_{k}\leftarrow\texttt{getIntersections}(\mathcal{B}^{\downarrow w},\Omega_{\varphi},k)$
[…]
@@ 11    $k\leftarrow k-1$
[…]
@@
The proof rests on the fact that, in the worst case, the parameter $k$ will reach a value of $1$. This will then correspond to the most-probable world of $\varphi$, which entails all possible monolithic explanations for $\varphi$. From Lemma 2, we know that the most-probable world of $\varphi$ entails all possible monolithic explanations for $\varphi$, that is, for any $\tilde{E}(\varphi)$, $\omega_{1}\models\tilde{\epsilon}$, and $\tilde{\epsilon}\subseteq I_{1}$. Therefore, as Algorithm 3 uses $I_{1}$ as an input to Algorithm 1, which is guaranteed to return a solution, the algorithm is also guaranteed to terminate with a solution. ∎
@@
First, note that if the user-defined parameter is initialized to $\hat{k}=|\Omega(\varphi)|$, then Algorithm 3 will perform an exhaustive and iterative search, starting from $k=|\Omega(\varphi)|$, to find $I_{k}$, such that $I_{k}\models\varphi$, and use it in Algorithm 1. Now, as the algorithm discounts $k$ by 1 at each new iteration, eventually it will be the case that $I_{k}\models\varphi$ and $I_{k+1}\not\models\varphi$. From Corollary 2, we then know that $\exists\tilde{\epsilon}\subseteq I_{k}$ such that $P(\tilde{\epsilon})\geq P(I_{k})\geq\overset{k}{\underset{i=1}{{{\sum}}}}P(\omega_{i})$, which means that $\tilde{\epsilon}$ corresponds to a $k$-bounded monolithic explanation. Therefore, the algorithm is guaranteed to return a maximal $k$-bounded monolithic explanation for $\varphi$. ∎
@@ 6.2.2 Model Reconciling Explanations
We now move on to the case of computing probabilistic model reconciling explanations $\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for an explanandum $\varphi$ from an agent knowledge base $\mathrm{KB}_{\alpha}$ for a human belief base $\mathcal{B}_{h}$. Similarly to what we described in Section 6.2.1, Algorithm 2 can directly be used on $\mathrm{KB}_{\alpha}$ and $\mathcal{B}_{h}^{\downarrow w}$ for computing model reconciling explanations. Additionally, the concept of a $k$-bounded explanation (Definition 20) can also be used to guarantee a lower bound on the probability of $\tilde{\epsilon}^{+}$.
@@
Result: A probabilistic model reconciling explanation $\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$
@@ 14    $k\leftarrow k-1$
[…]
@@ 7.1 Experimental Setup
Experiments were conducted on a system equipped with an M1 Max processor and 32GB of memory. The algorithms were implemented in Python, utilizing the PySAT toolkit (?) for SAT solving, MCS/MUS finding, weighted MaxSAT, and minimal hitting set computations. The time limit for all experiments was set to 500s500s. [^17] Code repository: https://github.com/YODA-Lab/Probabilistic-Monolithic-Model-Reconciling-Explanations.
@@
Note that we created associated belief bases for each problem by simply adding a random weight to each formula in the knowledge base.
@@ 7.2.1 Monolithic Explanations
Table 3: Number of Instances Solved (S) vs. Timed Out (T/O) by alg1 ($\hat{k}=1$) and alg3 ($\hat{k}=50$, $\hat{k}=100$, $\hat{k}=150$, $\hat{k}=200$). Note that Runtime denotes the average runtime over all solved instances.
@@
Table 3 tabulates the instances solved (i.e., found a monolithic explanation within the time limit) and not solved (i.e., timed out) by alg1 ($\hat{k}=1$) and alg3 at $\hat{k}=\{5,100,150,200\}$.191919alg3 at $\hat{k}=1$ corresponds to alg1 because each encoded knowledge base is consistent and entails the explanandum. As such, all formulae in the knowledge base are true in the most-probable world of the explanandum (i.e., $\hat{k}=1$), which means that alg3 reduces to alg1. We observe that the algorithm managed to solve most instances across different values of $\hat{k}$. Figure 3 shows the runtime distributions of alg3 across all values of $\hat{k}$ for computing a monolithic explanation. Interestingly, we observe that the runtimes decrease as $\hat{k}$ increases. This can be explained by the fact that for larger values of $\hat{k}$, alg3 considers the intersections of more worlds where the explanandum is true, which means that the number of formulae that are true in these intersections decreases. As such, the overall search space of monolithic explanations decreases as well, thus resulting in a reduced runtime needed for alg1 to compute a monolithic explanation. This can also be observed more granularly in Figure 4, where we can see the runtime distributions of alg1 ($\hat{k}=1$) and alg3 at $\hat{k}=200$ for each instance of the planning, scheduling, and random CNF problems. Again, the runtime of alg3 at $\hat{k}=200$ is smaller than that of alg1. Moreover, and as expected, in Figure 5, we can observe a positive correlation between runtime and the size of the encoded knowledge bases—as the size of the knowledge base increases, the runtimes increase as well. This is due to the fact that there is an increasing number of variables and formulae that must be considered, thus increasing the computational effort needed by the WMaxSAT, MCS, and hitting set solvers.
@@ 7.2.2 Model Reconciling Explanations
Table 4: Instances Solved (S) vs. Timed Out (T/O) for the Planning, Scheduling, and Random CNF Benchmarks for alg2 and alg4 at $\hat{k}=200$. Note that Runtime denotes the average runtime over all solved instances.
@@
Table 4 tabulates the instances solved and timed out by alg2 and alg4 at $\hat{k}=200$ across the five scenarios, where we observe the following trends. For the planning instances, the runtime of both algorithms increases as the difference between the models of the agent and human increases (Scenarios 1 to 5), since both algorithms search over the explanation search space, which increases as the number of differences between the two models increases. As in the previous experiments, alg4 at $\hat{k}=200$ yields faster runtimes than alg2. For the scheduling instances, we observe that the runtimes increase from Scenario 1 to 3, but decrease from Scenarios 4 to 5. Upon closer inspection, this is mainly because the instances solved in these scenarios were easier (i.e., smaller knowledge base sizes) than those solved in the other three scenarios, thus resulting in smaller average runtimes. A similar trend is observed for the random CNF instances. However, in the random CNF instances, alg2 managed to solve more instances than alg4. After examining them more closely, we found that the main bottleneck of alg4 in those instances was computing the most-probable worlds of the explanandum (i.e., the WMaxSAT solver). Even for smaller values of $\hat{k}$, the solver failed to compute all the worlds under the specified time limit—the increase in search space (e.g., because of considering $\mathcal{B}_{\alpha}$ and $\mathcal{B}_{h}$) increased the complexity of these instances. We expect that an optimized and more dedicated solver may be able to overcome this limitation. The runtime distributions for alg2 and alg4 across all commonly solved instances and across commonly solved instances in each scenario can be seen in Figures 6 and 7, respectively. For these instances, we observe, like in the previous experiments, that alg4 has faster runtimes than alg2.
Moreover, in Figure 8 we see the distributions of the model reconciling explanation lengths computed by both algorithms. As expected, the general trend is that the size of the explanation $\epsilon^{+}$ (i.e., formulae from $\mathrm{KB}_{\alpha}$ for $\mathrm{KB}_{h}$ (or $\mathcal{B}_{h}$)) increases with each scenario, as the difference between the agent and human models increase. The same trend can be seen for $\epsilon^{-}$—each scenario from 1 to 5 has an increasing amount of inconsistencies between the two models. Interestingly, $\epsilon^{-}$ was largest in the random CNF instances. This indicates that the inconsistencies between the human model and the corresponding $\epsilon^{+}$ were high. That can also be used to explain why alg4 failed to solve a subset of random CNF instances—highly inconsistent knowledge bases are considered as the most difficult instances for MaxSAT solvers.
@@ 8.1 Related Planning Work
We have briefly discussed the body of work within explainable AI planning (XAIP) on the topic of model reconciliation in Section 1 and situated our work within that literature. [^20] For a more complete discussion on XAIP related work, please see Section 9 of ? (?). For model reconciling explanations, a similar line of research that considers uncertainty about the human user’s model is that of ? (?). In their work, they focus on scenarios where the human’s model is located within a space of possible human models that the agent has, and their method operates in that space to find conformant explanations, i.e., explanations applicable to a set of possible models. In contrast, in our approach we represent an uncertain human model as a probability distribution that the agent has, and we have explicitly defined the notion of probabilistic explanation as well as metrics to measure its quality.
For monolithic explanations, most XAIP works have focused on contrastive explanations in deterministic settings. These explanations take the form of “Why not A (instead of B)?”, where A is an alternative (or foil) suggested by the human to a decision B proposed by the agent. Contrastive explanations have found applications in linear temporal logic systems (?), general epistemic accounts (?), multi-agent optimization problems (?), and in oversubscription planning (?). There have also been used to provide a taxonomy of user questions that often arise during interactive plan exploration (?), as well as towards creating user interfaces for decision support systems (?, ?).
@@ 8.2 Related KR Work
In the monolithic case, our definition of a probabilistic explanation (Definition 10) may appear similar to what was proposed by ? (?). However, an important distinction is that ? is dealing with epistemic states that do not contain the explanandum, while we are dealing with belief bases that do contain the explanandum. We also define a different notion of explanatory power as well as present algorithms for computing explanations. ? (?) have also considered the problem of defining what constitutes an explanation in probabilistic systems, however they solely focus on epistemic states defined over causal structures.
@@
The notion of (monolithic) explanation has also been explored by the probabilistic logic programming (PLP) community (?, ?), a formalism that extends logic programming languages with probabilities. In PLP, explanations have been associated with possible worlds. [^21] Recall that a possible world is a truth value assignment to the atoms in the language. The most prominent task there is that of the most probable explanation (MPE), which consists of finding the world with the highest probability given some evidence (?). However, a world does not show the chain of inferences of a given explanandum and it is not minimal by definition, since it usually includes a (possibly large) number of probabilistic facts whose truth value is irrelevant for the explanandum. An alternative approach is using the proof of an explanandum as an explanation (?), where a proof is a (minimal) partial world $\omega^{\prime}$ such that for all worlds $\omega\supseteq\omega^{\prime}$, the explanandum is true in $\omega$. In this case, one can easily ensure minimality, but even if the partial world contains no irrelevant facts, it is still not easy to determine the chain of inferences behind a given explanandum. Finally, ? (?) have leveraged explanations in PLP as approximation techniques for more efficiently computing weighted model counting problems.
It is important to mention that logic-based methods have also been employed in machine learning problems (?). For example, ? (?) proposed CLIME, a constraint-driven explanation framework for black-box ML models that enables users to specify Boolean constraints to guide the perturbation phase when generating explanations. Unlike their approach that focuses on constraining the input space during perturbation, our approach integrates uncertainty directly into the explanation generation process through probabilistic logic. Both approaches use concepts from formal methods, but our work addresses the specific case of explanations in probabilistic scenarios and model reconciliation, providing a complementary perspective to CLIME’s constraint-driven framework. A similar line of research is the work by ? (?), where they introduced an approach to computing probabilistic explanations across various classifier types (such as decision trees and naive Bayes classifiers). Their framework addresses the challenge of generating minimal sets of features that guarantee a prediction with a certain probability threshold. While ? focus on efficiently computing explanations for specific classifier architectures, our framework provides a more general theoretical foundation for explanation generation in uncertain environments and extends to the model reconciliation problem.
@@
In the model reconciliation case, we have extended our previous work on the logic-based model reconciliation problem (?) to handle scenarios where the human model is uncertain. To the best of our knowledge, the application of probabilistic explanations in the context of model reconciliation that we consider in this work is novel.
@@
Finally, the algorithms presented in this paper are an extension of our previous work (?). Specifically, Algorithms 1 and 2 are inspired by a procedure for computing a smallest minimal unsatisfiable set (SMUS) of an inconsistent formula, originally presented by ? (?). The method is also related to other similar approaches for enumerating MUSes and minimal correction sets (MCSes). Moreover, our approach is similar in spirit to the HS-tree presented by ? (?). Although the original purpose was to enumerate diagnoses, Reiter’s procedure can be easily adapted to enumerate MUSes (called conflicts in that paper) as already noted by ? (?). However, the computation of an SMUS might require more substantial modifications. Procedures like the one presented by Reiter, which target MCSes (diagnoses) instead of MUSes (conflicts), can be seen as the dual version of our algorithm. In particular, the algorithm MaxHS (?) applies the same idea of iteratively computing and testing a minimal hitting set for the computation of a MaxSAT solution (the complement of the smallest MCSes).
There are other approaches that exploit the duality between MUSes and MCSes, but instead of iteratively checking if the current hitting set is an MUS, they first compute the set of all MCSes (?). This has the potential advantage that once all the MCSes are known, every minimal hitting set on the collection of all MCSes is guaranteed to be an MUS (Proposition 1). However, as the number of MCSes is, in the worst case, exponential in the size of the formula, this approach might fail even before reporting the first MUS. This is particularly unnecessary when the target is to return a single support, like the one presented in this paper.
9 Discussion
In this paper, we presented a logic-based framework for generating probabilistic explanations within uncertain knowledge bases (e.g., belief bases). We distinguished between two notions of explanation—monolithic, where explanations are defined with respect to an agent knowledge base, and model reconciling, where explanations are defined with respect to an agent knowledge base as well as a human user knowledge base. Particularly, we introduced formal definitions for probabilistic monolithic and probabilistic model reconciling explanations, along with metrics to evaluate their quality (explanatory gain and explanatory power).
Our framework operates under several assumptions that we need to address. First, a fundamental assumption is that the problem domain can be encoded in a logical language. This assumption holds for many structured decision-making scenarios but may not apply to domains where knowledge is primarily subsymbolic or where the underlying decision processes are difficult to formalize in a logical language. The general effectiveness of our approach thus depends critically on how well the domain can be represented within a logical framework. Further, we used propositional (probabilistic) logic as our language of choice due to its simplicity, capacity to encode a plethora of domains, as well as computationally efficiency (e.g., through SAT solvers). While our methods and algorithms are presented within propositional logic, their underlying principles are broadly applicable to any constraint system where the satisfiability of subsets can be decided. This opens avenues for extending our work to other logical systems, such as Markov Logic Networks (?) and Probabilistic Logic Programs (?).
For model reconciling explanations specifically, our framework makes some assumptions about human inferential capabilities. Notably, we presume that human users possess the reasoning capacity to process and understand the explanations provided by the AI agent. However, this does not imply that humans can reason with the same efficiency as AI agents—much like how humans can perform arithmetic operations correctly but are significantly slower than calculators. Rather, we assume that given sufficient time and a properly formulated explanation, humans can validate the logical correctness of the AI agent’s decision-making process.
Another big assumption in model reconciliation is that the agent has direct access to the human’s belief base. This assumption, while necessary for the operationalizing our framework, is often unrealistic in practice. Thus, a promising direction for future research involves approximating the human model through iterative interactions. Indeed, in a recent work (?), we showed how (propositional) probabilistic logic can be effectively used to learn a human model based on past, dialogue-based interactions. This line of work provides a fruitful ground for our framework to create more personalized and effective model reconciliation processes.
At the other end of the spectrum, a practical challenge of explainable decision-making frameworks involves how the explanations are communicated to human users. [^22] Note that, in certain settings, it has been shown that explanations in the form of model reconciliation are well understood and preferred by human users (?, ?, ?, ?, ?). While we tackled the problem of generating explanations with formal guarantees about their quality, we did not directly explore the presentation of these explanations to users. There are several avenues to explore here. First, the emergence of large language models (LLMs) (?) may offer an immediate solution to this problem. Specifically, one can use LLMs as translation engines between logic and natural language, and thus create hybrid systems that generate natural language explanations with formal guarantees (provided by logical framework). For example, we presented a system for tackling this idea in the domain of coursework planning (?, ?). In future work, we plan to extend it for general problems expressible in propositional logic. It is also worth emphasizing that explanations can also be communicated in other forms, such as visualizations, something that has been shown to improve the users’ understanding when compared to text-based explanations alone (?, ?).
Finally, it is important to acknowledge the resurgence of hybrid approaches that combine symbolic and neural-based methods (?, ?). These neuro-symbolic AI systems aim to synergize the strengths of both paradigms: the robust learning capabilities and pattern recognition of neural networks with the interpretability and reasoning power of symbolic systems. We position the methods proposed in this work as complementary to the advancements in neuro-symbolic AI, potentially informing and enhancing the explainability aspects of future neuro-symbolic systems.